An organic compound had the following composition 37.21% carbon, 7.75% hydrogen and rest chlorine . Determine the molecular formula of the compound given that the molecular compound is 65(c=12H 1 CL=35.5)
1
Expert's answer
2020-12-14T14:50:15-0500

Solutions:

parameters:

molar mass of the compound

Carbon = 37.21%.

Hydrogen = 7.75%.

Chlorine = 100% - (37.21% + 7.75%)

= 100% - (44.96%)

= 55.04%.

C H Cl

% of elements= 37.21 7.75 55.04


mass in (g)= 37.21 7.75 55.04


mole= 37.21 / 12 7.75 / 1 55.04 / 35.5

mole = 3.10. 7.75 1.55


Divide through.

by the smallest value.= 3.10 / 1.55 7.55 / 1.55 1.55 / 1.55


mole ratio = 2 5 1


The emprical formula of the compound is C2H5Cl.

Recall that M ef × n = molar mass of the compound.

M ef = mass of the emprical formula.

n = number of repeating units of the emprical formula present in one mole of the molecule.

(C2H5Cl)n = 65

[ (12 × 5) + (1 × 5) + (35.5 × 1) ] n = 65.

(24 + 5 + 35.5) = 65

(64.5)n = 65.

n = 65 / 64.5

n = 1.

(C2H5Cl) × 1 = C2H5Cl

So, the molecular formula of the compound is C2H5Cl